Let us consider a swing as a harmonic oscillator! The velocity of the swing at the lowest position is 3 m/s. The acceleration at the highest point is 3 m/s2 . Write down the displacement-time function!


1
Expert's answer
2021-06-21T11:38:27-0400

"v=A\omega=3,"

"a=A\omega^2=3," "\implies"

"\omega=1~\frac{rad}{s},"

"A=3~m," "\implies"

"x=A\sin \omega t,"

"x=3\sin t~(m)."